f a R200 billion increases in investment spending creates R200 billion of new
income in the first round of the multiplier process and R160 billion in the second
round. Calculate:
a. the marginal propensity to consume (MPC).
b. the value of the expenditure multiplier in this closed economy.
1
Expert's answer
2020-04-21T18:39:26-0400
  1. a"\Delta Y(1-MPC)=\Delta I"

200(1-MPC)=200

MPC=0

b."M=\frac{\Delta Y}{\Delta I}=\frac{200}{200}=1"


2 a."\Delta Y(1-MPC)=\Delta I"

160 =(1-MPC)=200

MPC=0.25

b."M=\frac{\Delta Y}{\Delta I}=\frac{160}{200}=0.8"